OpenStudy (anonymous):

what is the value of sin(x+π/3)?

OpenStudy (bittuaryan):

\[\sin(\frac{ \pi }{ 3 }+x)=\sin(\frac{ \pi }{ 3 })cosx+\cos(\frac{ \pi }{ 3 })+sinx=\frac{ \sqrt{3}}{ 2 }cosx+\frac{ 1}{ 2 }sinx\]

OpenStudy (anonymous):

you are right

OpenStudy (bittuaryan):

Thanks @Meinme

OpenStudy (bittuaryan):

Medal!..wow....thanks @leehk9402

OpenStudy (anonymous):

no problem :D @bittuaryan